You are here: Foswiki>Commissioning Web>Logs>17Jun2012_Log (18 Jun 2012, JuanGuerra)Edit Attach

AO/PISCES . INAF SCIENCE NIGHTS

Observer: DMiller (Tucson), JGuerra, KBoutsia (LBTO)
Telescope Operator: SAllanson (LBTO)
Additional AO Support: RBriguglio, CArcidiacono (LBTO) APuglisi
Telescope Support: JUrban(LBTO)
Instrument Support: DOfficer (LBTO)

Summary

Telescope configure to LBC, the detector red was repaired. (fiber problem) Half night lost due to storm. @11:45 switch to PISCES+AO. seeing variable and clouds passing forced to open the loop.

one hour after midnight the sky started to be more stable and the loop is more stable.

Details

-The first target is a science target a nebulosity targtet and the loop couldn't be close due to clouds passing, the clouds forced the loop open.

-The loop is close in a bright source 9 and 12 magnitude without any proble and the system is getting stable working with brighter reference star.

The loop is broken after taking a sky images for about 10 minutes, the loop was in pause and he tried to resume the target was off the piramid and the loop is open. Preset was failed and it was necessary to send again the preset on-axis.

At the end of the night the bcu39 failed and this produced the preset to fail. The bcu39 was recovered and the preset was sent again.

-- JuanGuerra - 17 Jun 2012
Topic revision: r3 - 18 Jun 2012, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback